OpenStudy (anonymous):

cos(x + pi/6) + sin(x - pi/3) = A ; what is A?

OpenStudy (anonymous):

sin(α – β) = sin(α)cos(β) – cos(α)sin(β) cos(α + β) = cos(α)cos(β) – sin(α)sin(β)

OpenStudy (anonymous):

use these formulae's accordingly and let me know what your answer is.

OpenStudy (anonymous):

these are angle-sum and difference identities.